# How weak are weak factors? Uniform inference for signal strength in signal plus noise models

### v1

@@ 8.1. The function $\mathcal{G}(w)$
The first sum is finite, and, therefore, almost surely converges to 00 at speed $|w|^{-2}$ as $|w|\to\infty$. The second sum has at most $|w|^{3/2}$ terms and each term is upper-bounded as $O(|w|^{-2})$; hence, the sum is $O(w^{-1/2})$. The terms of the last sum can be upper-bounded by a constant times $j^{2\varepsilon-4/3}$, and therefore the sum is upper bounded by a constant times $|w|^{\frac{3}{2}(2\varepsilon-1/3)}$. Combining all three bounds, we arrive at (8.26).
For the second sum in (8.25), we do summation by parts:
$\sum_{j=1}^{M}\frac{\xi_{j}^{2}-1}{w+\left(\frac{3\pi j}{2}\right)^{2/3}}=\frac{1}{w+\left(\frac{3\pi(M+1)}{2}\right)^{2/3}}\sum_{k=1}^{M}(\xi_{k}^{2}-1)\\ -\sum_{j=1}^{M}\left[\sum_{k=1}^{j}[\xi_{k}^{2}-1]\right]\left(\frac{1}{w+\left(\frac{3\pi(j+1)}{2}\right)^{2/3}}-\frac{1}{w+\left(\frac{3\pi j}{2}\right)^{2/3}}\right).$
@@
The last expression clearly tends to 00 as $|w|\to\infty$, uniformly in $M$. We conclude that for large $w$, up to $o(1)$ error only the first term in (8.25) contributes to the $w\to\infty$ asymptotics. This term is deterministic and we can analyze it by converting sums into integrals. For large $x$ and large $w$ with $\mathrm{Re}(w)\geq 0$, we have
@@
We conclude that
(8.29) $\begin{split}&\lim_{x\to-\infty}\left[\left(\sum_{j:\,\left(\frac{3\pi j}{2}\right)^{2/3}<-x}\frac{1}{w+\left(\frac{3\pi j}{2}\right)^{2/3}}\right)-\frac{2}{\pi}\sqrt{-x}\right]\\ &\qquad=-\frac{2}{\pi}\sqrt{w}\lim_{x\to-\infty}\arctan\left(\frac{\sqrt{-x}}{\sqrt{w}}\right)+O(|w|^{-1/2})\\ &\qquad=-\frac{2}{\pi}\sqrt{w}\lim_{x\to-\infty}\left(\frac{\pi}{2}-\arctan\left(\frac{\sqrt{w}}{\sqrt{-x}}\right)\right)+O(|w|^{-1/2})=-\sqrt{w}+O(|w|^{-1/2}).\end{split}$
@@
In order to prove (8.3), we again use (8.25). (8.26) and (LABEL:eq_x58) imply that the sum of the first and the third terms is $-\sqrt{w}+o\left(w^{-1/4}\right)$ and it remains to analyze the second term. Note that it is a sum of mean 00 independent random variables, and therefore the Central Limit Theorem applies. Hence, it remains to compute the asymptotic variance of the sum as $w\to\infty$, which is
(8.30) $\sum_{j=1}^{\infty}\frac{2}{\left(w+\left(\frac{3\pi j}{2}\right)^{2/3}\right)^{2}}=(2+o(1))\int\limits_{0}^{\infty}\frac{\mathrm{d}x}{\left(w+\left(\frac{3\pi x}{2}\right)^{2/3}\right)^{2}}=(2+o(1))\frac{1}{2\sqrt{w}},$
which matches the claim of (8.2). ∎
@@
Using Definition 4.2, the main computation of the proof is to replace $\mathcal{G}(w)$ with $-\Theta$ and then solve (8.31), viewed as an equation on unknown $w$, and treating $\Theta$ as a parameter. In this way we get the desired equivalent form of (8.4):
@@
In order to justify the validity of this computation, we use the monotonicity of $\mathcal{G}(w)$ on $[\mathfrak{a}_{1},+\infty)$. The distributional limit of $\frac{\mathcal{T}(\Theta)-\Theta^{2}}{2\sqrt{\Theta}}$ is obtained from the computation of the following probabilities for $t\in\mathbb{R}$, in which we used Definition 4.2:
@@ 8.2. A class of meromorphic functions
In this subsection we first consider deterministic meromorphic complex functions. We then introduce randomness and establish general theorems that aid in proving convergence to $\mathcal{G}(w)$.
@@
[…]
@@
For $R>1$, the magnitude of the sum is bounded by (8.32) and each term goes to 00 as $R\to\infty$. Hence, by the dominated convergence theorem the sum goes to 00 and $f(\mathbf{i}R)-f(\mathbf{i})=o(R)$. Since obviously also $f(\mathbf{i})=o(R)$ as $R\to\infty$, the conclusion follows. ∎
@@
and there exists a deterministic function ϕ:+:italic-$\phi:\mathbb{R}_{+}\to\mathbb{C}$ such that for any $\varepsilon>0$ and $R>0$
@@
[…]
Theorem 8.13 is inspired by Aizenman and Warzel (2015), see Theorem 3.1 and Section 6 there, as well as Sodin (2018, Section 1.4). There are, however, important differences: in Aizenman and Warzel (2015) the function $\phi(R)$ was constant – this does not hold in our main application. Additionally, our topology of convergence is stronger than that of Aizenman and Warzel (2015), which will be crucial for us when we want to solve equations of the form $f_{n}(z)=\theta$. Unlike Sodin (2018), we do not make use of any results from complex analysis, beyond the basic theory.
@@
Let $f_{n}$, $n=1,2,\dots$, and $f$ be deterministic functions from $\Omega_{-}$, corresponding to $(\gamma^{n},\{x_{j;n}\}_{j=1}^{\infty},\{w_{j;n}\}_{j=1}^{\infty})$, and $(\gamma,\{x_{j}\}_{j=1}^{\infty},\{w_{j}\}_{j=1}^{\infty})$ Suppose that:
@@
[…]
@@
The first factor stays uniformly bounded as $n\to\infty$, and it remains to show that the second factor tends to 00 as $M\to\infty$ (uniformly in $n$). For that we observe
@@
which is a well-defined finite random variable due to $f\in\Omega_{-}$ and (8.32). In order to justify the interchange of the order of summation and taking the $n\to\infty$ limit, we need to produce an additional tail bound. We observe that for large $Q-q$ we have
@@
[…]
@@
In Theorem 8.13, instead of $f_{n},f\in\Omega_{-}$, we can take $f_{n}=f^{+}_{n}+f^{-}_{n}$, $f=f^{+}+f^{-}$, with $f_{n}^{\pm},f^{\pm}\in\Omega_{\pm}$ corresponding to $(\gamma_{n}^{\pm},\{x_{j;n^{\pm}}\}_{j=1}^{\infty},\{w_{j;n}^{\pm}\}_{j=1}^{\infty})$ and $(\gamma^{\pm},\{x_{j}^{\pm}\}_{j=1}^{\infty},\{w_{j}^{\pm}\}_{j=1}^{\infty})$. Replacing (8.34) with
@@
[…]
@@
For both $f\in\Omega_{-}$ and $f\in\Omega_{+}$, the function $-f(z)$ belongs to the Herglotz-Nevanlinna class, meaning it maps the upper half-plane to itself. Our proofs can be adapted to handle most other Herglotz–Nevanlinna functions, as long as they exhibit sublinear growth. However, the linear Herglotz-Nevanlinna function $az$, $a>0$, is excluded from our definitions: Lemma 8.12 is essential for the arguments.
@@ 8.3. Convergence to $\mathcal{G}(w)$
In this section we use Theorem 8.13 to derive sufficient conditions for convergence towards $\mathcal{G}(w)$ of Theorem 8.1. For each $N=1,2,\dots$, we are given an $N$–tuple of random variables $\lambda_{1;N}\geq\lambda_{2;N}\geq\dots\geq\lambda_{N;N}$; we assume that for each $N$ these variables are defined on its own probability space, which therefore depends on $N$. In addition, for each $N$ we are given a sequence of i.i.d. Gaussian random variables $\xi_{1},\xi_{2},\dots,\xi_{N}$ on the same $N$–dependent probability space (we omit an additional $N$ from the index of $\xi_{i}$, because the distribution of $\xi_{i}$ does not depend on it), which are independent of $\{\lambda_{i;N}\}_{i=1}^{N}$. In addition, we are given a continuous non-negative function $h(x)\geq 0$. We encode the data via the empirical Stieltjes transform:
@@
where $\mu(x)\mathrm{d}x$ is an independent of $N$ probability measure $\mu(x)\mathrm{d}x$, supported on a finite interval $[\lambda_{-},\lambda_{+}]\subset\mathbb{R}$, and which is a limit of the empirical measures of $\lambda_{i;N}$.
@@
We denote
(8.47) $\mathcal{G}_{N}(w)=N^{1/3}\mathfrak{s}^{-2/3}\left(\frac{1}{N}\sum_{j=1}^{N}\frac{h(\lambda_{j;N})\,\xi_{j}^{2}}{\lambda_{+}+N^{-2/3}\mathfrak{s}^{-2/3}w-\lambda_{j;N}}-\mathfrak{m}\right).$
The next theorem asserts that one can couple the processes defined for the different values of $N$ in such a way that $\mathcal{G}_{N}$ converge to the Airy-Green function.
@@
Under Assumption 8.19, there exists a coupling that places random variables ($\bigl{(}\{\lambda_{j;N}\}_{j=1}^{N},\{\xi_{j}\}_{j=1}^{N}\bigr{)}_{N=1}^{\infty}$ and the Airy-Green function $\mathcal{G}(w)$ on the same probability space, such that almost surely
[…]
@@
For any finitely many deterministic $w_{1},\dots,w_{k}\in\mathbb{C}$, the convergence in (8.48) holds in joint distribution over $w=w_{1}$, …, $w=w_{k}$.
@@
Fix $\theta\in\mathbb{R}$ and let $\tilde{w}_{N}$ denote the largest real solution of the equation $\mathcal{G}_{N}(w)=\theta$ and let $\tilde{w}$ denote the largest real solution of the equation $\mathcal{G}(w)=\theta$. Under Assumption 8.19,
@@
For some settings we might have more than one sequence $\xi_{j}$ and more than one choice for $h(x)$. One can take finite $K$ and consider simultaneous edge limits for
@@
Condition 1. We need to check that the functions $\mathcal{G}_{N}$ and $\mathcal{G}$ are almost surely in the class $\Omega_{-}$. For $\mathcal{G}_{N}$ this is immediate. $\mathcal{G}$ also belongs to $\Omega_{-}$, as can be seen by rewriting the definition:
@@
Condition 3. It remains to check (8.35). We set $\phi(R)=-\frac{1+\mathbf{i}}{\sqrt{2}}\sqrt{R}$. Proposition 8.3 yields that $\mathcal{G}(\mathbf{i}R)+\frac{1+\mathbf{i}}{\sqrt{2}}\sqrt{R}$ goes to 00 as $R\to\infty$ almost surely, and, therefore, also in probability, verifying the first limit in (8.35). For the second limit we prove that
@@
By the Markov inequality, (8.50) is sufficient for establishing (8.35). The limit (8.50) is a corollary of the assumptions (8.45) and (8.46). Indeed, taking the expectation with respect to $\xi_{j}$ first and using $\mathbb{E}(\xi_{j}^{2}-1)^{2}=2$, we have
@@
Directly subtracting the values at two points inside the definition of $\mathcal{G}(w)$, we see that it is a strictly monotone decreasing function of real $w>\mathfrak{a}_{1}$. Proposition 8.3 implies that it varies from +$\infty$+ ∞ to $\infty$- ∞ on $(\mathfrak{a}_{1},+\infty)$. We conclude that for each $\theta\in\mathbb{R}$, there exists a unique random $\tilde{w}>\mathfrak{a}_{1}$, such that $\mathcal{G}(\tilde{w})=\theta$.
Next, we consider the coupling of Theorem 8.20 and note that by Rouché’s theorem (see e.g., Ahlfors (1979)) locally-uniform convergence of meromorphic functions implies the convergence of their zeros. Hence, $\tilde{w}_{N}\to\tilde{w}$ almost surely, and, therefore, also in distribution. ∎
@@
In order to verify the conditions of Theorem 8.20, we are going to use the following universal bound valid for all $\beta$–ensembles. Take $\beta>0$, a function $V(x):[a,b]\to\mathbb{R}$, often called a potential, and consider for each $N=1,2,\dots$, a probability measure on $\{\lambda_{i}\}_{i=1}^{N}$, such that $b\geq\lambda_{1}>\lambda_{2}>\dots>\lambda_{N}\geq a$, with probability density proportional to
@@
[…]
@@
In all three cases $[\lambda_{-},\lambda_{+}]$ denotes the support of $\mu_{V}$. Let us also introduce the empirical and limiting Stieltjes transforms:
@@
Note a deterministic bound $\left|\frac{1}{z-\lambda_{i}}\right|\leq\frac{1}{\eta}$. This bound allows us to discard (for the purpose of proving (8.55)) various events, whose probabilities are decaying fast with $N$. We will do this several times in the proof.
Let us first assume that $\lambda_{-}>0$ (which happens whenever $p$ is bounded away from 00). By the large deviations principle for the support of the empirical measure of eigenvalues (see, e.g., Borot and Guionnet (2013, Proposition 2.1)) for each $\delta>0$, with probability exponentially close to $1$ for large $N$ all the eigenvalues $\lambda_{1},\dots,\lambda_{N}$ lie in the positive segment $[\lambda_{-}-\delta,\lambda_{+}+\delta]\subset(0,+\infty)$. We choose a contour $\Gamma$ enclosing both this segment and a point $z$, intersecting real axis in positive points, and lying inside the domain of validity of (8.54). We write using (8.53) and residue expansion of the contour integral:
@@
For the first term in (8.58), the integrand is upper-bounded by a deterministic constant, because $\Gamma_{1}$ is away from $[\lambda_{-}-\delta,\lambda_{+}+\delta]$ by construction and we ignore the event when eigenvalues are not in this segment, as explained at the beginning of the proof. The length of $\Gamma_{1}$ is of order $1/N$, and therefore the first term is upper-bounded by $C_{1}N^{-2}$.
@@
In the case $\lambda_{-}=0$ we need to be more careful, because if we argue in the same way as for $\lambda_{-}>0$, then $\Gamma$ has to loop around 00 and $\sqrt{u}$ is no longer holomorphic on the contour. Instead, we choose the contour $\Gamma$ passing directly through 00 and enclosing $(0,\lambda_{+}+\delta)$. Compared to the previous argument, it is no longer true that on the part of the contour at distance $\frac{1}{N}$ from 0 the integrand is upper-bounded by a constant. However, in this case the splitting of the contour $\Gamma$ is not required at all: the $1/\eta$ singularity of the bound (8.54) is compensated by $\sqrt{u}$ factor in the integrand, and the resulting expression $\eta^{-1/2}$ is integrable at $\eta=0$. ∎
@@
For JOE ensembles, the bound (8.54) further extends to the cases $h(x)=\sqrt{1-x}$ and $h(x)=\sqrt{x(1-x)}$, in the same form: there exist $\tilde{\eta}>0$ and $C>0$, such that for all $N\geq 1$, and $z=E+\mathbf{i}\eta$ with $0<\eta<\tilde{\eta}$, $\max(\lambda_{-}-\tilde{\eta},0)<E<\min(\lambda_{+}+\tilde{\eta},1)$, we have
@@
We only consider the case $h(x)=1$, as other cases are similar. We would like to upper bound the absolute value of the difference,
$\frac{1}{N}\sum_{i=1}^{N}\frac{1}{z-\lambda_{i;N}}-\frac{1}{N}\sum_{i=1}^{N}\frac{1}{z-\mu_{i;N}}.$
The real part of the difference is

### v2

@@ 8.1. The function $\mathcal{G}(w)$
The first sum is finite, and, therefore, almost surely converges to 0 at speed $|w|^{-2}$ as $|w|\to\infty$. The second sum has at most $|w|^{3/2}$ terms and each term is upper-bounded as $O(|w|^{-2})$; hence, the sum is $O(w^{-1/2})$. The terms of the last sum can be upper-bounded by a constant times $j^{2\varepsilon-4/3}$, and therefore the sum is upper bounded by a constant times $|w|^{\frac{3}{2}(2\varepsilon-1/3)}$. Combining all three bounds, we arrive at (8.25).
For the second sum in (8.24), summation by parts converts it into:
$\frac{1}{w+\left(\frac{3\pi(M+1)}{2}\right)^{2/3}}\sum_{k=1}^{M}(\xi_{k}^{2}-1)\\ -\sum_{j=1}^{M}\left[\sum_{k=1}^{j}[\xi_{k}^{2}-1]\right]\left(\frac{1}{w+\left(\frac{3\pi(j+1)}{2}\right)^{2/3}}-\frac{1}{w+\left(\frac{3\pi j}{2}\right)^{2/3}}\right).$
@@
The last expression tends to 0 as $|w|\to\infty$, uniformly in $M$. Hence, for large $w$, up to $o(1)$ error only the first term in (8.24) contributes to the $w\to\infty$ asymptotics. This term is deterministic and approximates an integral. For large $x$ and $w$ with $\mathrm{Re}(w)\geq 0$, we have
@@
We conclude that the first term in (8.24) is asymptotically
(8.28) $\frac{2}{\pi}\sqrt{w}\lim_{x\to-\infty}\arctan\left(\frac{\sqrt{-x}}{\sqrt{w}}\right)+O(|w|^{-1/2})\\ =-\frac{2}{\pi}\sqrt{w}\lim_{x\to-\infty}\left(\frac{\pi}{2}-\arctan\left(\frac{\sqrt{w}}{\sqrt{-x}}\right)\right)+O(|w|^{-1/2})=-\sqrt{w}+O(|w|^{-1/2}).$
@@
In order to prove (8.3), we again use (8.24). (8.25) and (8.28) imply that the sum of the first and the third terms is $\sqrt{w}+o\left(w^{-1/4}\right)$ and it remains to analyze the second term. It is a sum of mean 0 independent random variables, and the Central Limit Theorem applies. Hence, it remains to compute the asymptotic variance of the sum as $w\to\infty$, which is
(8.29) $\sum_{j=1}^{\infty}\frac{2}{\left(w+\left(\frac{3\pi j}{2}\right)^{2/3}\right)^{2}}=(2+o(1))\int\limits_{0}^{\infty}\frac{\mathrm{d}x}{\left(w+\left(\frac{3\pi x}{2}\right)^{2/3}\right)^{2}}=(2+o(1))\frac{1}{2\sqrt{w}}.\qed$
@@
Using Definition 4.2, the main computation of the proof is to replace $\mathcal{G}(w)$ with $\Theta$ and then solve (8.30), viewed as an equation on unknown $w$, and treating $\Theta$ as a parameter. In this way we get the desired equivalent form of (8.4):
@@
In order to justify the validity of this computation, we use the monotonicity of $\mathcal{G}(w)$ on $\mathfrak{a}_{1},+\infty)$. The distributional limit of $\frac{\mathcal{T}(\Theta)-\Theta^{2}}{2\sqrt{\Theta}}$ is obtained from the computation of the following probabilities for $t\in\mathbb{R}$, in which we used Definition 4.2:
@@ 8.2. A class of meromorphic functions
We first consider deterministic functions, then introduce randomness and establish theorems that aid in proving convergence to $\mathcal{G}(w)$.
@@
The sum (8.32) converges uniformly on any compact subset of $\mathbb{C}\setminus\{x_{j}\}_{j=1}^{\infty}$; that is, it converges in the topology $\overset{\text{mer}}{\rightarrow}$.
@@
For $R>1$, the magnitude of the sum is bounded by (8.31) and each term goes to 0 as $R\to\infty$. Hence, by the dominated convergence theorem the sum goes to 0 and $f(\mathbf{i}R)-f(\mathbf{i})=o(R)$. Since obviously also $f(\mathbf{i})=o(R)$ as $R\to\infty$, the conclusion follows. ∎
@@
and there exists a deterministic function $\phi:\mathbb{R}_{+}\to\mathbb{C}$ such that for any $\varepsilon>0$ and $R>0$
@@
Then there exists a coupling that places all $f_{n}$ and $f$ on the same probability space such that almost surely $f_{n}\overset{\text{mer}}{\rightarrow}f$.
As noted earlier, this implies convergence in distribution $f_{n}\overset{\text{mer, d}}{\rightarrow}f$, and thus convergence in distribution of $(f_{n}(z_{j}))_{j=1}^{k}$ to $(f(z_{j}))_{j=1}^{k}$ at any finite set of points.
Theorem 8.13 is inspired by Aizenman and Warzel (2015), see Theorem 3.1 and Section 6 there, as well as Sodin (2018, Section 1.4). There are, however, important differences: in Aizenman and Warzel (2015) the function $\phi(R)$ was constant – this does not hold in our main application. Additionally, our topology of convergence is stronger than that of Aizenman and Warzel (2015), which is crucial when solving equations of the form $f_{n}(z)=\theta$. Unlike Sodin (2018), we do not make use of any results from complex analysis, beyond the basics.
@@
Let $f_{n}$, $n=1,2,\dots$, and $f$ be deterministic functions from $\Omega_{-}$, corresponding to $(\gamma^{n},\{x_{j;n}\}_{j=1}^{\infty},\{w_{j;n}\}_{j=1}^{\infty})$, and $(\gamma,\{x_{j}\}_{j=1}^{\infty},\{w_{j}\}_{j=1}^{\infty})$ . Suppose that:
@@
Then $f_{n}\overset{\text{mer}}{\rightarrow}f$, i.e., $f_{n}(z)\to f(z)$, uniformly over $z$ in compact subsets of $\mathbb{C}\setminus\{x_{j}\}_{j=1}^{\infty}$.
@@
The first factor stays uniformly bounded as $n\to\infty$, and it remains to show that the second factor tends to 0 as $M\to\infty$ (uniformly in $n$). For that we observe
@@
which is a well-defined finite random variable due to $f\in\Omega_{-}$ and (8.31). In order to justify the interchange of the order of summation and taking the $n\to\infty$ limit, we need to produce an additional tail bound. We observe that for large QqQ-q we have
@@
By the Skorokhod’s representation theorem (see, e.g., Billingsley (1999, Chapter 1, Theorem 6.7)) one can construct a coupling, so that (8.41) becomes almost sure convergence. Then Lemma 8.15 implies $f_{n}\overset{\text{mer}}{\rightarrow}f$. ∎
@@
In Theorem 8.13, instead of $f_{n},f\in\Omega_{-}$, we can take $_{n}=f^{+}_{n}+f^{-}_{n}, f=f$$f^{+}+f^{-}$, with $f_{n}^{\pm},f^{\pm}\in\Omega_{\pm}$ corresponding to $(\gamma_{n}^{\pm},\{x_{j;n^{\pm}}\}_{j=1}^{\infty},\{w_{j;n}^{\pm}\}_{j=1}^{\infty})$ and $(\gamma^{\pm},\{x_{j}^{\pm}\}_{j=1}^{\infty},\{w_{j}^{\pm}\}_{j=1}^{\infty})$. Replacing (8.33) with
@@
and keeping (8.34), the conclusion of Theorem 8.13 continues to hold: almost surely $f_{n}\overset{\text{mer}}{\rightarrow}f$.
@@
For $f\in\Omega_{+}$, the function f(z)-f(z) lies in the Herglotz-Nevanlinna class, mapping the upper half-plane to itself. Our proofs extend to other sublinear Herglotz-Nevanlinna functions but exclude the linear $az$, $a>0$, since Lemma 8.12 is essential.
@@ 8.3. Convergence to $\mathcal{G}(w)$
In this section we use Theorem 8.13 to derive sufficient conditions for convergence towards $\mathcal{G}(w)$ of Theorem 8.1. For each $N=1,2,\dots$, we are given an NN–tuple of random variables $\lambda_{1;N}\geq\lambda_{2;N}\geq\dots\geq\lambda_{N;N}$; we assume that for each $N$ these variables are defined on its own probability space, which therefore depends on $N$. In addition, for each $N$ we are given a sequence of i.i.d. Gaussian random variables $\xi_{1},\xi_{2},\dots,\xi_{N}$ on the same NN–dependent probability space (we omit an additional $N$ from the index of $\xi_{i}$, because the distribution of $\xi_{i}$ does not depend on it), which are independent of $\{\lambda_{i;N}\}_{i=1}^{N}$. In addition, we are given a continuous non-negative function $h(x)\geq 0$. The data is assumed to satisfy the following condition, using the empirical Stieltjes transform notation:
@@
where $\mu(x)\mathrm{d}x$ is an independent of $N$ probability measure $\mu(x)\mathrm{d}x$, supported on a finite interval $\lambda_{-},\lambda_{+}]\subset\mathbb{R}$, and which is a limit of the empirical measures of $\lambda_{i;N}$.
@@
Under Assumption 8.19 and denoting
(8.46) $\mathcal{G}_{N}(w)=N^{1/3}\mathfrak{s}^{-2/3}\left(\frac{1}{N}\sum_{j=1}^{N}\frac{h(\lambda_{j;N})\,\xi_{j}^{2}}{\lambda_{+}+N^{-2/3}\mathfrak{s}^{-2/3}w-\lambda_{j;N}}-\mathfrak{m}\right),$
there exists a coupling that places random variables $\bigl(\{\lambda_{j;N}\}_{j=1}^{N},\{\xi_{j}\}_{j=1}^{N}\bigr)_{N=1}^{\infty}$ and the Airy-Green function $\mathcal{G}(w)$ on the same probability space, such that almost surely
(8.47) $\mathcal{G}_{N}(w)\overset{\text{mer}}{\rightarrow}h(\lambda_{+})G(w),\qquad N\to\infty.$
@@
For any finitely many deterministic $w_{1},\dots,w_{k}\in\mathbb{C}$, the convergence in (8.47) holds in joint distribution over $w_{1}, …, w=w$$w_{k}$.
@@
For $\theta\in\mathbb{R}$, let $\tilde{w}_{N}$ denote the largest real solution of the equation $\mathcal{G}_{N}(w)=\theta$ and let $\tilde{w}$ denote the largest real solution of the equation $\mathcal{G}(w)=\theta$. Under Assumption 8.19,
@@
We might have more than one sequence $\xi_{j}$ and more than one $h(x)$. One can take finite $K$ and consider simultaneous edge limits for
@@
Condition 1. We need to check that the functions $\mathcal{G}_{N}$ and $\mathcal{G}$ are in the class $\Omega_{-}$. For $\mathcal{G}_{N}$ this is immediate. $\mathcal{G}$ also belongs to $\Omega_{-}$, as can be seen by rewriting the definition:
@@
Condition 3. It remains to check (8.34). We set $\phi(R)=-\frac{1+\mathbf{i}}{\sqrt{2}}\sqrt{R}$. Proposition 8.3 yields that $\mathcal{G}(\mathbf{i}R)+\frac{1+\mathbf{i}}{\sqrt{2}}\sqrt{R}$ goes to 0 as $R\to\infty$ almost surely, and, therefore, also in probability, verifying the first limit in (8.34). For the second limit we prove that
@@
By the Markov inequality, (8.49) is sufficient for establishing (8.34). Taking the expectation with respect to $\xi_{j}$ first and using $\mathbb{E}(\xi_{j}^{2}-1)^{2}=2$, we have
@@
Directly subtracting the values at two points inside the definition of $\mathcal{G}(w)$, we see that it is a strictly monotone decreasing function of real $w>\mathfrak{a}_{1}$. Proposition 8.3 implies that it varies from $\infty$ to -$\infty$ on $(\mathfrak{a}_{1},+\infty)$. We conclude that for each $\theta\in\mathbb{R}$, there exists a unique random $\tilde{w}>\mathfrak{a}_{1}$, such that $\mathcal{G}(\tilde{w})=\theta$.
Next, we consider the coupling of Theorem 8.20. By Rouché’s theorem (see e.g., Ahlfors (1979)) locally-uniform convergence of meromorphic functions implies the convergence of their zeros. Hence, $\tilde{w}_{N}\to\tilde{w}$ almost surely, and, therefore, also in distribution. ∎
@@
To verify the conditions of Theorem 8.20, we use the following universal bound valid for all $\beta$–ensembles. Take $\beta>0$, a function $V(x):[a,b]\to\mathbb{R}$, called a potential, and consider for each $N=1,2,\dots$, a probability measure on $\{\lambda_{i}\}_{i=1}^{N}$, such that $b\geq\lambda_{1}>\lambda_{2}>\dots>\lambda_{N}\geq a$, with probability density proportional to
@@
  • For GOE, $\mu_{V}$ is the semicirle law of density $\frac{1}{2\pi}\sqrt{4-x^{2}}\,\mathbf{1}_{[-2,2]}\,\mathrm{d}x$.
@@
In all three cases $\lambda_{-},\lambda_{+}]$ denotes the support of $\mu_{V}$. Let us also introduce the empirical and limiting Stieltjes transforms:
@@
Note a deterministic bound $\left|\frac{1}{z-\lambda_{i}}\right|\leq\frac{1}{\eta}$. This bound allows us to discard events, whose probabilities are decaying fast with $N$. We will do this several times.
Let us first assume that $\lambda_{-}>0$ (which happens whenever $p$ is bounded away from 0). By the large deviations principle for the support of the empirical measure of eigenvalues (see, e.g., Borot and Guionnet (2013, Proposition 2.1)) for each $\delta>0$, with probability exponentially close to $1$ for large $N$ all the eigenvalues $\lambda_{1},\dots,\lambda_{N}$ lie in the positive segment $\lambda_{-}-\delta,\lambda_{+}+\delta]\subset(0,+\infty)$. We choose a contour $\Gamma$ enclosing both this segment and a point $z$, intersecting real axis in positive points, and lying inside the domain of validity of (8.53). We write using (8.52) and residue expansion of the contour integral:
@@
For the first term in (8.57), the integrand is upper-bounded by a deterministic constant, because $\Gamma_{1}$ is away from $\lambda_{-}-\delta,\lambda_{+}+\delta]$ by construction and we ignore the event when eigenvalues are not in this segment, as explained at the beginning of the proof. The length of $\Gamma_{1}$ is of order $1/N$, and therefore the first term is upper-bounded by $C_{1}N^{-2}$.
@@
In the case $\lambda_{-}=0$ we need to be more careful, because if we argue in the same way as for $\lambda_{-}>0$, then $\Gamma$ has to loop around 0 and $\sqrt{u}$ is no longer holomorphic on the contour. Instead, we choose the contour $\Gamma$ passing directly through 0 and enclosing $(0,\lambda_{+}+\delta)$. Compared to the previous argument, it is no longer true that on the part of the contour at distance $\frac{1}{N}$ from 0 the integrand is upper-bounded by a constant. However, in this case the splitting of the contour $\Gamma$ is not required at all: the $1/\eta$ singularity of the bound (8.53) is compensated by $\sqrt{u}$ factor in the integrand, and the resulting expression $\eta^{-1/2}$ is integrable at $\eta=0$. ∎
@@
For JOE ensembles, the bound (8.53) further extends to the cases $h(x)=\sqrt{1-x}$ and $h(x)=\sqrt{x(1-x)}$: there exist $\tilde{\eta}>0$ and $C>0$, such that for all $N\geq 1$, and $z=E+\mathbf{i}\eta$ with $0<\eta<\tilde{\eta}$, $\max(\lambda_{-}-\tilde{\eta},0)<E<\min(\lambda_{+}+\tilde{\eta},1)$, we have
@@
We only consider the case $h(x)=1$, as other cases are similar. We would like to upper bound the absolute value of the difference, $\frac{1}{N}\sum_{i=1}^{N}\frac{1}{z-\lambda_{i;N}}-\frac{1}{N}\sum_{i=1}^{N}\frac{1}{z-\mu_{i;N}}.$ The real part of the difference is